myininaya (myininaya):

Evaluate the following limit: \[\lim_{n \rightarrow \infty}(\sum_{i=1}^{4,000}i^\frac{1}{n}-3,999)^n\]

ganeshie8 (ganeshie8):

im trying to mess with below limit bu the sum looks complicated to simplify.. \[\large e^x = \lim \limits_{n\to \infty} \left(1 + \dfrac{x}{n}\right)^n\]

myininaya (myininaya):

\[\lim_{n \rightarrow \infty}(\sum_{i=1}^{4,000}i^\frac{1}{n}-3,999)^n\] Notice we can write the above as: \[\lim_{n \rightarrow \infty}((1^\frac{1}{n}-1)+(2^\frac{1}{n}-1)+(3^\frac{1}{n}-1)+ \cdots + (4000^\frac{1}{n}-1)+1)^n\] this a^z-1 thing should look kind of familiar \[\text{ Let } T(x)=a^x\] \[\text{ So } T'(x)=a^x \ln(a) =\lim_{z \rightarrow x} \frac{a^z-a^x}{z-x} \] \[\text{ Then } T'(0)=ln(a)=\lim_{z \rightarrow 0} \frac{a^z-1}{z}\] Ok do you see that a^z -1 we can use that for our problem if we let z=1/n so \[\ln(a)=\lim_{n \rightarrow \infty } \frac{a^\frac{1}{n}-1}{\frac{1}{n}} \] For large n there exist small epsilon such that \[\ln(a)=\frac{a^\frac{1}{n}-1}{\frac{1}{n}} + \epsilon \] This still holds if we multiply both sides by 1/n \[\frac{1}{n} \ln(a)=a^\frac{1}{n}-1+ \frac{1}{n} \epsilon \] since 1/n*epsilon approaches 0 for large n we can use the approximation: \[\frac{1}{n} \ln(a) \approx a^\frac{1}{n}-1\] So let's look at the thing we rewrote earlier : \[\lim_{n \rightarrow \infty}((1^\frac{1}{n}-1)+(2^\frac{1}{n}-1)+(3^\frac{1}{n}-1)+ \cdots + (4000^\frac{1}{n}-1)+1)^n\] \[\lim_{n \rightarrow \infty}(\frac{1}{n} \ln(1)+\frac{1}{n} \ln(2)+\frac{1}{n} \ln(3)+ \cdots +\frac{1}{n} \ln(4000)+1)^n \] \[\lim_{n \rightarrow \infty}(\frac{1}{n} \ln(4000!)+1)^n \] \[\text{ Let } y=(\frac{\ln(4000!)}{n}+1)^n \] \[\lim_{n \rightarrow \infty} \ln(y)=\lim_{n \rightarrow \infty} \frac{\ln(1+\frac{\ln(4000!)}{n})}{\frac{1}{n} }=\lim_{n \rightarrow \infty} \frac{\frac{\frac{-\ln(4000!)}{n^2}}{1+\frac{\ln(4000!)}{n}}}{-\frac{1}{n^2}} \\ =\lim_{n \rightarrow \infty}\frac{\frac{-\ln(4000!)}{n^2+\ln(4000!)n}}{\frac{-1}{n^2}} \\ =\lim_{n \rightarrow \infty}\frac{\ln(4000!)n^2}{n^2+\ln(4000!)n}=\ln(4000!)\] Therefore \[\lim_{n \rightarrow \infty}(\sum_{i=1}^{4,000}i^\frac{1}{n}-3,999)^n\] = \[e^{\ln(4000!)}=4000!\]

myininaya (myininaya):

yeah that is the exact formula I used

myininaya (myininaya):

I used an approximation.

myininaya (myininaya):

If anyone has any other way of looking at this let me know.

myininaya (myininaya):

I usually play with small numbers first.

myininaya (myininaya):

you know that have this form

ganeshie8 (ganeshie8):

beautiful !! i wouldn't have gotten this clever idea : \(\large \frac{1}{n} \ln(a) \approx a^\frac{1}{n}-1\) i think the number 3999 exists exactly for this purpose xD

myininaya (myininaya):

I remember a very similar problem to this I had a lot of trouble with it I think it was on the Putnam exam 7 years ago (maybe more years ago) It is weird how some things you don't get until way later

myininaya (myininaya):

\[\lim_{n \rightarrow \infty}(\sum_{i=1}^{k}i^\frac{1}{n}-(k-1))^n=k!\]

myininaya (myininaya):

@ganeshie8 why did you think that one formula would be involved? just curious. what was your thinking?

myininaya (myininaya):

was it simply because of the ( )^n that you thought it

ganeshie8 (ganeshie8):

Oh, i have worked a lot on my notes - since the n is in the exponent, outside the sum, I thought of simplifying the sum first before taking the limit.... i got stuck at that part... once the sum simplifies, rest of the problem is trivial

ganeshie8 (ganeshie8):

lol exactly ^

myininaya (myininaya):

i think you did a lot better than me seeing this problem for the first time when i seen it i was so completely lost at first

OpenStudy (ikram002p):

follow *

OpenStudy (thomas5267):

WOW! This is ridiculous man!

OpenStudy (ikram002p):

i cant think of anything better that was amazing approach !